# Consumption Partial Insurance

### v4

@@ 7 Conclusion
We measure the degree of consumption partial insurance to income shocks, accounting for higher-order moments of the income distribution. Such moments, in particular left skewness and excess kurtosis, are important features of unexplained income growth, documented in many recent papers. We use recent data from the PSID and establish several new findings.
@@
Second, in estimating a linear consumption model over the recent years, we find a low pass-through of permanent shocks (large insurance) and full insurance to transitory ones. We attribute one third of the difference from BPP to the consumption imputation from the CEX, another third to the biennial nature of our data, and the remainder to time effects.
@@
Fourth, in estimating a nonlinear consumption function, we find that bad permanent shocks transmit more strongly than good ones, and the pass-through increases with the severity of the shock. By contrast, transitory shocks are fully insured in the full sample, but they are not insured among the less wealthy. The nature of income risk and consumption transmission vary by age and the household’s position in the income distribution.
Fifth, tail risk incurs substantial welfare costs, up to 1/8 of lifetime consumption.
@@ Appendix A Derivation of the consumption function
In this appendix we derive the consumption function under a permanent-transitory specification for log income. We follow similar derivations in Blundell et al. (2008), BPP hereafter, though our approach is a refinement of theirs given the second-order approximation we carry out herein. We show below how to derive the most general (quadratic) consumption function (5); the linear function (4) is a special case of this. We let a time unit in the model correspond to one calendar year. However, data in the PSID in recent years are available only every second year, so we adapt the derivations to reflect this.
@@ B.3 Quadratic consumption function parameters
where the matrix of coefficients $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks. [^1] With biennial data, matrix $\mathbf{A}$ is given by (different from $\mathbf{A}$ with annual data in the paper) $\mathbf{A}$ $=\left(\begin{array}[]{lllll}2\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ 2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2}))&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&4\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&0\\ 0&-\gamma_{v_{t}}&-2\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&-\kappa_{v_{t}}-\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}&-2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ The determinant of this matrix is given by
@@
How do different levels of skewness and kurtosis connect with the transmission parameters? Consider the transmission of permanent shocks, abstracting from transitory shocks (i.e., $\psi^{(1)}_{t}=\psi^{(2)}_{t}=\omega^{(22)}_{t}=0$). $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are pinned down by $2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}+2\gamma_{\zeta_{t}}\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})$, which corresponds to the first equation in (B.1). Absent any transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and permanent shocks. Suppose $\phi^{(1)}_{t}$ is fixed. We highlight three cases:
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}>0$. Consumption growth and permanent shocks co-vary more strongly than in the linear consumption process, i.e., there is excess co-movement. For $\gamma_{\zeta_{t}}<0$, this configuration identifies $\phi^{(2)}_{t}<0$ and implies that negative shocks transmit more strongly than other shocks. Negative shocks are often the largest shocks (due to left skewness), which explains the excess co-movement. Ceteris paribus, a less skewed distribution of shocks ($\gamma_{\zeta_{t}}$ drops in magnitude) implies a more negative $\phi^{(2)}_{t}$, because a more symmetric distribution associated with fewer large shocks requires a stronger transmission of said shocks to maintain a given excess co-movement. If $\gamma_{\zeta_{t}}>0$, this configuration identifies $\phi^{(2)}_{t}>0$ and analogous arguments apply to positive shocks.
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}<0$. Consumption growth and permanent shocks co-move less strongly than in the linear process, i.e., there is excess smoothness. For $\gamma_{\zeta_{t}}<0$, this identifies $\phi^{(2)}_{t}>0$ and suggests a muted pass-through of negative shocks. These are the largest shocks to income, so muting their pass-through suppresses the covariance between consumption and income growth. The opposite is true if $\gamma_{\zeta_{t}}>0$.
@@
In practice, $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are jointly pinned down by this and other equations of system (B.1), and they are also determined jointly with the transmission parameters of transitory shocks, to which analogous statements apply – skipped here for brevity. The sensitivity analysis is thus more nuanced so simulations can prove useful. Starting from the baseline estimates of the income process (table 2), reducing the left skewness of permanent shocks by half increases the magnitude of $\phi^{(2)}_{t}$ by almost 10% (more negative). Independently reducing the kurtosis of permanent shocks by half increases the magnitude of $\phi^{(2)}_{t}$ by 60% (more negative). The response of $\phi^{(1)}_{t}$ is much smaller, confirming that the values of skewness and kurtosis are crucial for identifying the pass-through of tail shocks.
An interesting special case arises when the distributions of permanent and transitory shocks are normal, have the same variance between them, and there is no measurement error. In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}$, $\phi_{t}^{(1)}=(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)+(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)+(1/(6\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+2}\right)$, where $\sigma_{t}$ is the dispersion of the generic distribution. Even in such a special environment, $\phi_{t}^{(2)}$, the transmission parameter of tail shocks, is separately identified from $\phi_{t}^{(1)}$. Therefore, identification does not require the shocks to be non-Gaussian, but skewness and kurtosis (both defined also in the Gaussian case) are still crucial for identifying the pass-through of tail shocks. Despite normality, shocks in the (normal) tails are associated with a different pass-through from that of average shocks. Analogous expressions separately identify $\psi_{t}^{(1)}$ and $\psi_{t}^{(2)}$.
@@ C.3 Sources of tail income shocks
Table C.3 provides evidence on the sources of large income reductions that contribute to income left skewness and excess kurtosis. Columns 1 & 2 report average unemployment, disability, bad health, etc, in the bottom 10% and top 90% of the distribution of residual income changes $\Delta y_{it}$, respectively. Column 3 reports the $p$-value of the difference between the two; in most cases, the bottom 10% (effectively households facing large income reductions) experience unemployment, worsening health, large drops in hours, and precarious housing at much higher rates than the rest.
@@ C.4 CEX data and imputation in the PSID
Sample selection and variables. To parallel the design of the PSID sample, we use CEX interview data between 1999 and 2019. We select a sample of households that mimics closely our baseline selection in the PSID. [^3] To improve the quality of the food demand estimation subsequently, we keep all calendar years between 1999 and 2019, i.e., not just the odd ones as in the PSID. Specifically, we select continuously married couples with the male spouse aged 30 to 65. We require non-missing data on expenditure and basic demographics, and we drop those with zero food expenditure given the nature of the imputation subsequently. Similar to the PSID, we define consumption as the sum of real expenditure on nondurable goods and services, namely food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses, education, and daycare. As the CEX collects quarterly data over four rolling quarters, we generate annual consumption as the sum of consumption over four quarters. We assign the final observation to the calendar year to which the underlying quarterly data mostly correspond to. Our final sample consists of 31,751 observations.
Table C.4 presents a comparison of means between the PSID and CEX samples for a host of household characteristics over the period 1999–2019. The two samples are very close with respect to age, family size, number of children, race, education, region of residence, or labor market participation. Household earnings are higher in the PSID than in the CEX; BPP noted this also for the earlier period (1980–1992) and argued that it is due to the more comprehensive definition of income in the PSID. Food expenditure in the PSID is higher, on average, by about 60% than food expenditure in the CEX. This is in contrast to BPP who find that food expenditure is similar across the two surveys over 1980–1992. This results in consumption being about 25% higher in the PSID than in the CEX.
@@
C.5 Comparison using the original BPP data
@@
First, we replicate BPP over the original data/period of time, using BPP’s annual growth rates of income and consumption. This is column 1 in table C.5. We obtain numerically similar results to BPP, namely $\phi^{(1)}=0.64$ while $\psi^{(1)}$ is statistically zero.
Second, we calculate biennial growth rates of income and consumption and re-estimate the linear model. This is column 2 in table C.5; this estimation is analogous to the linear function estimation in column 1 of table 3, albeit using the original BPP sample. Note that despite the use of biennial data, $\phi^{(1)}$ and $\psi^{(1)}$ still express the pass-through at annual rates (see appendix B). $\phi^{(1)}$ drops by 20% to $0.518$, while $\psi^{(1)}$ remains qualitatively unaffected. With biennial data, we only identify the pass-though of an ‘aggregate’ permanent shock over two years. Some higher frequency shocks are muted at the observed lower frequency, given that we can only observe an average of two one-year permanent shocks over the two-year period. In other words, the use of biennial data allows us to only measure an ‘aggregate’ or average pass-through within a given biennial period, and some higher-frequency shocks are inevitably smoothed out at this observed lower frequency.
@@
There are some caveats comparing the exercise in table C.5 and our results with the most recent data in table 3. First, our main exercise uses consumption data internally available in the PSID, while table C.5 uses consumption imputed from the CEX. We showed earlier that the imputation imparts substantial imputation error. Second, while the estimation in table C.5 does not correct the income moments for measurement error (in line with BPP), our main exercise in table 3 does so (in line with Blundell et al., 2016). Finally, the use of biennial growth rates does not allow the identification of $\theta$, the moving average parameter of the transitory shock, so the transitory shock is $MA(0)$ in our baseline ($\theta=0$), while it is $MA(1)$ in the original BPP exercise.
@@
Notes: The table presents the estimates of the parameters of the linear consumption function, using the original BPP data over 1980–1992 (PSID income data, consumption data imputed from the CEX). The data are available through Blundell et al. (2008)’s online replication package. Columns 1-2 allow the target second-order income and consumption moments to vary with calendar time, while columns 3-4 target said moments imposing time-invariance. Columns 1 and 3 estimate the model using annual growth rates of income and consumption, while columns 2 and 4 use biennial rates. In line with BPP, estimation is done via diagonally weighted GMM; GMM standard errors are in parentheses. The very small unimportant numerical differences between column 1 herein and the original BPP are likely due to different software and optimization algorithms (we use Matlab and fmincon).

### v5

@@ 6 Conclusion
We introduce a simple analytical framework to measure the degree of consumption partial insurance to income shocks, accounting for higher-order moments of the income distribution. Such moments, in particular left skewness and excess kurtosis, are important features of unexplained income growth, documented recently in the U.S. and other countries.
We use recent income and consumption data from the PSID and establish several findings.
@@
Second, in estimating a linear consumption function over the recent years, we find a low pass-through of permanent shocks (large insurance) and full insurance to transitory ones. We attribute one third of the difference from BPP to the consumption imputation from the CEX, another third to the biennial nature of our data, and the remainder to time effects.
@@
Fourth, in estimating a nonlinear consumption function, we find that bad permanent shocks transmit more strongly than good ones, and the pass-through increases with the severity of the shock. By contrast, transitory shocks are mostly fully insured. The nature of income risk and consumption transmission vary substantially by age and the household’s position in the income distribution.
@@ Appendix A Derivation of the consumption function
This appendix shows how to derive the consumption function under a permanent-transitory specification for log income. We follow similar derivations in Blundell et al. (2008), BPP hereafter, though our approach is a refinement of theirs given the second-order approximation we carry out herein. We show below how to derive the most general (quadratic) consumption function (5); the linear function (4) is a special case of this. We let a time unit in the model correspond to one calendar year. However, data in the PSID in recent years are available only every second year, so we adapt the derivations to reflect this.
@@ B.3 Quadratic consumption function parameters
where the matrix of coefficients $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks. [^1] With biennial data, matrix $\mathbf{A}$ is different from $\mathbf{A}$ with annual data in the main paper and is given by $\mathbf{A}$ $=\left(\begin{array}[]{lllll}2\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ 2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2}))&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&4\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&0\\ 0&-\gamma_{v_{t}}&-2\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&-\kappa_{v_{t}}-\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}&-2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ The determinant of this matrix is given by
@@
How do different levels of skewness and kurtosis influence the transmission parameters? Consider the transmission of permanent shocks, abstracting from transitory shocks, i.e., $\psi^{(1)}_{t}=\psi^{(2)}_{t}=\omega^{(22)}_{t}=0$. $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are pinned down by the first equation in (B.1), namely $2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}+2\gamma_{\zeta_{t}}\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})$. Absent any transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and permanent shocks. Suppose $\phi^{(1)}_{t}$ is fixed. We highlight three cases:
@@
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}>0$. Consumption growth and permanent shocks co-vary more strongly than the linear consumption process suggests, i.e., there is excess co-movement. For $\gamma_{\zeta_{t}}<0$, this configuration identifies $\phi^{(2)}_{t}<0$ and implies that negative shocks transmit more strongly than other shocks. Negative shocks are often the largest shocks (due to left skewness), which explains the excess co-movement. Ceteris paribus, a less skewed distribution ($\gamma_{\zeta_{t}}$ drops in magnitude) implies a more negative $\phi^{(2)}_{t}$, because a more symmetric distribution associated with fewer large shocks requires a stronger transmission of said shocks to maintain a given excess co-movement. If $\gamma_{\zeta_{t}}>0$, this configuration identifies $\phi^{(2)}_{t}>0$ and analogous arguments apply to positive shocks.
  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}<0$. Consumption growth and permanent shocks co-move less strongly than the linear process suggests, i.e., there is excess smoothness. For $\gamma_{\zeta_{t}}<0$, this identifies $\phi^{(2)}_{t}>0$ and suggests a muted pass-through of negative shocks. These are the largest shocks to income, so muting their pass-through suppresses the covariance between consumption and income growth. The opposite is true if $\gamma_{\zeta_{t}}>0$.
@@
In practice, $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are jointly pinned down by these and other equations of system (B.1), and they are also determined jointly with the transmission parameters of transitory shocks, to which analogous statements apply – skipped here for brevity. The sensitivity analysis is thus more nuanced, so simulations can prove useful. Starting from the baseline estimates of the income process (table 2), reducing the left skewness of permanent shocks by half increases the magnitude of $\phi^{(2)}_{t}$ by almost 10% (more negative). Independently reducing the kurtosis of permanent shocks by half increases the magnitude of $\phi^{(2)}_{t}$ by 60% (more negative). The response of $\phi^{(1)}_{t}$ is much smaller, confirming that the values of skewness and kurtosis are crucial for identifying the pass-through of tail shocks.
An interesting special case arises when the distributions of permanent and transitory shocks are normal, have the same variance between them, and there is no measurement error. In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}$, $\phi_{t}^{(1)}=(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)+(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)+(1/(6\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+2}\right)$, where $\sigma_{t}$ is the dispersion of the hypothetical distribution. Even in such a special environment, $\phi_{t}^{(2)}$, the transmission parameter of tail shocks, is separately identified from $\phi_{t}^{(1)}$. Identification does not require the shocks to be non-Gaussian, but skewness and kurtosis (both defined also in the Gaussian case) are still crucial for identifying the pass-through of tail shocks. Despite normality, shocks in the (normal) tails are associated with a different pass-through from that of average shocks. Analogous expressions separately identify $\psi_{t}^{(1)}$ and $\psi_{t}^{(2)}$.
@@ C.3 Sources of tail income shocks
Table C.3 presents evidence on the underlying sources of large income reductions that drive the left skewness and excess kurtosis of income growth. Columns 1 & 2 report average unemployment, disability, bad health, etc, in the bottom 10% and top 90% of the distribution of residual income changes $\Delta y_{it}$, respectively. Column 3 reports the $p$-value of the difference between the two; in most cases, the bottom 10% (effectively households facing large income reductions) experience unemployment, worsening health, large drops in hours, and precarious housing at much higher rates than the rest.
@@ C.4 CEX data and imputation in the PSID
Sample selection and variables. To parallel the design of the PSID sample, we use CEX interview data between 1999 and 2019. We select a sample of households that mimics closely our baseline selection in the PSID. [^3] To improve the quality of the food demand estimation subsequently, we keep all calendar years between 1999 and 2019, i.e., not just the odd ones as in the PSID. Specifically, we select continuously married couples with the male spouse aged 30 to 65. We require non-missing data on expenditure and basic demographics, and we drop those with zero food expenditure given the nature of the imputation subsequently. As in the PSID, we define consumption as the sum of real expenditure on nondurable goods and services, namely food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses, education, and daycare. As the CEX collects quarterly data over four rolling quarters, we generate annual consumption as the sum of consumption over four quarters. We assign the final observation to the calendar year to which the underlying quarterly data mostly correspond to. Our final sample consists of 31,751 observations.
Table C.4 presents a comparison of means between the PSID and CEX samples for a host of household characteristics over the period 1999–2019. The two samples are very close with respect to age, family size, number of children, race, education, region of residence, and labor market participation. Household earnings are higher in the PSID than in the CEX; BPP noted this also for the earlier period (1980–1992) and argued that it is due to the more comprehensive definition of income in the PSID. Food expenditure in the PSID is higher, on average, by about 60% than food expenditure in the CEX. This is in contrast to BPP who find that food expenditure is similar across the two surveys over 1980–1992. This results in consumption being about 25% higher in the PSID than in the CEX.
@@
C.5 Estimation using the original BPP data
@@
First, we replicate BPP over the original data/period of time, using our code but BPP’s annual growth rates of income and consumption. This is column 1 in table C.5. We obtain numerically similar results to BPP, namely $\phi^{(1)}=0.64$ while $\psi^{(1)}$ is statistically zero.
Second, we calculate biennial growth rates of income and consumption and re-estimate the linear model. This is column 2 in table C.5; this estimation is analogous to the linear function estimation in column 1 of table 3, albeit using the original BPP sample. Note that despite the use of biennial data, $\phi^{(1)}$ and $\psi^{(1)}$ still express the pass-through at annual rates (see appendix B). $\phi^{(1)}$ drops by 20% to $0.518$, while $\psi^{(1)}$ remains small but turns statistically significant. With biennial data, we only identify the pass-though of an ‘aggregate’ permanent shock over two years. Some higher frequency shocks are muted at the observed lower frequency, given that we can only observe an average of two one-year permanent shocks over the two-year period. In other words, the biennial data allow us to measure only an ‘aggregate’ or average pass-through within a given period, and some higher-frequency shocks are inevitably smoothed out at this observed lower frequency.
@@
There are some caveats comparing the exercise in table C.5 and our results from the recent PSID in table 3. First, our main exercise uses consumption data internally available in the PSID, while table C.5 uses consumption imputed from the CEX. We showed earlier that the imputation imparts substantial imputation error. Second, while the estimation in table C.5 does not correct the income moments for measurement error (in line with BPP), our main exercise in table 3 does so (in line with Blundell et al., 2016). Finally, the use of biennial growth rates does not allow the identification of $\theta$, the moving average parameter of the transitory shock, so the transitory shock is $MA(0)$ in our baseline ($\theta=0$), while it is $MA(1)$ in the original BPP exercise.
@@
Notes: The table presents the estimates of the parameters of the linear consumption function, using the original BPP data over 1980–1992 (PSID income data, consumption data imputed from the CEX). The data are available through Blundell et al. (2008)’s online replication package. Columns 1-2 allow the target second-order income and consumption moments to vary with calendar time, while columns 3-4 target said moments imposing time-invariance. Columns 1 and 3 estimate the model using annual growth rates of income and consumption, while columns 2 and 4 use biennial rates. In line with BPP, estimation is done via diagonally weighted GMM; GMM standard errors are in parentheses.
The small unimportant numerical differences between column 1 herein and the original BPP estimates are likely due to different software and optimization algorithms (our code uses Matlab and fmincon while BPP use Gauss).